### v1

# LLMs for Law: Evaluating Legal-Specific LLMs on Contract Understanding

Abstract
Despite advances in legal NLP, no comprehensive evaluation covering multiple legal-specific LLMs currently exists for contract classification tasks in contract understanding. To address this gap, we present an evaluation of 10 legal-specific LLMs on three English-language contract understanding tasks and compare them with 7 general-purpose LLMs. The results show that legal-specific LLMs consistently outperform general-purpose models, especially on tasks requiring nuanced legal understanding. Legal-BERT and Contracts-BERT establish new SOTAs on two of the three tasks, despite having 69% fewer parameters than the best-performing general-purpose LLM. We also identify CaseLaw-BERT and LexLM as strong additional baselines for contract understanding. Our results provide a holistic evaluation of legal-specific LLMs and will facilitate the development of more accurate contract understanding systems.
LLMs for Law: Evaluating Legal-Specific LLMs on Contract Understanding
Amrita Singhthanks: These authors contributed equally to this work., H. Suhan Karaca11footnotemark: 1, Aditya Joshi, Hye-young Paik, Jiaojiao Jiang School of Computer Science and Engineering University of New South Wales (UNSW), Sydney
1 Introduction
Recent work suggests that open-source legal-specific LLMs offer a promising, cost-effective, and privacy-preserving alternative to general-purpose LLMs Bhambhoria et al. (2024); Chalkidis et al. (2020).
Prior Work / Ours
Legal-Specific LLMs UNFAIR-ToS LEDGAR LEXDEMOD
Legal-BERT ✓/✓ ✓/✓ ✕/✓
Contracts-BERT ✕/✓ ✕/✓ ✓/✓
Legal-RoBERTa ✕/✓ ✕/✓ ✕/✓
CaseLaw-BERT ✓/✓ ✓/✓ ✕/✓
PoL-BERT ✕/✓ ✕/✓ ✕/✓
InLegalBERT ✕/✓ ✕/✓ ✕/✓
InCaseLawBERT ✕/✓ ✕/✓ ✕/✓
CustomInLawBERT ✕/✓ ✕/✓ ✕/✓
LexLM ✕/✓ ✕/✓ ✕/✓
Legal-XLM-R ✕/✓ ✕/✓ ✕/✓
Table 1: Comparison of our legal-specific LLMs evaluation and coverage with prior work across three contract datasets (tasks): ‘UNFAIR-ToS’ (Unfair Contractual Terms Identification, Lippi et al. (2019); Chalkidis et al. (2022)), ‘LEDGAR’ (Contract Provision Topic Classification, Tuggener et al. (2020); Chalkidis et al. (2022)), and ‘LEXDEMOD’ (Agent-Specific Deontic Modality Detection, Sancheti et al. (2022)). Terms in inverted commas refer to dataset names, tasks are in parentheses. ✓= model inclusion, ✕= model exclusion.
However, despite their advantages, these models remain significantly underutilized in current legal NLP downstream tasks. As illustrated in Table 1, legal-specific LLMs are rarely evaluated in prior work on three popular and freely available contract understanding tasks such as Unfair Contractual Terms Identification Lippi et al. (2019); Chalkidis et al. (2022), Contractual Provision Topic Classification Tuggener et al. (2020); Chalkidis et al. (2022), and Agent-Specific Deontic Modality Detection Sancheti et al. (2022). Despite the legal nature of documents/tasks, researchers have continued to favor general-purpose LLMs over legal-specific LLMs. In some cases, legal-specific LLMs are excluded entirely. For instance, recent studies such as Guha et al. (2023) and Singh et al. (2024), which explicitly focus on legal downstream tasks, do not include any legal-specific LLMs in their benchmarking evaluations. Therefore, this paper addresses the Research Question (RQ): How do legal-specific LLMs perform compared to general-purpose LLMs on nuanced legal tasks like contract understanding?
Dataset
Contract
Type
Task Task Type
Train/Dev/
Test Instances
Classes
UNFAIR-ToS
Chalkidis et al. (2022)
Terms of
Service
(Consumer
Contract)
Unfair Contractual
Terms Identification
Multi-label
Classification
5,532/2,275/
1,607
9
LEDGAR
Chalkidis et al. (2022)
Exhibit-10
Material
Contract
Contract Provision
Topic Classification
Multi-class
Classification
60,000/10,000/
10,000
100
LEXDEMOD
Sancheti et al. (2022)
Lease
Contract
Agent-Specific Deontic
Modality Detection
Multi-label
Classification
4,282/330/
1,777
7
Table 2: Overview of Datasets used for Benchmarking Legal-specific LLMs.
To address this question, we present a comprehensive evaluation of 10 open-source legal-specific LLMs with 7 general-purpose LLMs across the three distinct contract understanding tasks. Our results reveal consistent improvements in performance for legal-specific LLMs, particularly on tasks where legal and domain-specific semantics are critical. This benchmark serves as a resource for the community, offering a clearer understanding of model suitability and performance across tasks and model types. The contributions of this work are as follows: (a) To the best of our knowledge, we present the first benchmarking of multiple legal-specific LLMs across multiple contract understanding tasks; (b) We systematically compare their performance with that of general-purpose LLMs; (c) We identify model strengths, weaknesses, and task-specific challenges, offering insights for future research and deployment.
2 Contract Classification Tasks and Datasets
2.1 Dataset Selection Desiderata
Based on following factors, we select the legal contract datasets and tasks:
Language: English-language contract datasets are selected due to their availability, provide consistent benchmarking for future legal-specific models in the global research community, enable comparison with past benchmarked models.
Relevance and Diversity: The focus is on contract classification tasks that reflect real-world contract review and analysis challenges, and that test a model’s understanding of legal language, structure, and semantics. As shown in Table 2, three distinct tasks are selected, each using a different dataset and representing a unique contract classification scenario in terms of dataset size and number of classes.
Difficulty: Datasets are chosen where SOTA general-purpose language models do not achieve near-perfect performance Lippi et al. (2019); Tuggener et al. (2020); Sancheti et al. (2022), ensuring that benchmarking legal-specific language models remains challenging.
Availability & Size: Public, well-documented datasets are used, each large enough for stable training and evaluation. Proprietary, non-public, and very small datasets (under 3K sentences) are avoided to ensure reproducibility and generalizability. This criterion modifies and adapts the selection guidelines of Chalkidis et al. (2022).
Legal-Specific Model Pre-training Corpora # Doc Base Model
Legal-BERT
Chalkidis et al. (2020)
EU Legislation, UK Legislation, European Court of Justice (ECJ)
Cases, European Court of Human Right (ECHR) Cases, US Court
Cases, US Contracts
354K BERT-base-uncased
Contracts-BERT
Chalkidis et al. (2020)
US Contracts 76K BERT-base-uncased
Legal-RoBERTa
Geng et al. (2021)
Patent Litigations, US Court Cases, Google Patents Public Data
- RoBERTa-base
CaseLaw-BERT
Zheng et al. (2021)
Harvard Case Law (US federal and State courts)
3.4M BERT-base-uncased
PoL-BERT
Henderson et al. (2022)
Court Opinions, Government, Publications, Contracts, Statutes,
Legal Analyses, Regulations, and, more from US and EU
10M RoBERTa-large
InLegalBERT
Paul et al. (2023)
Indian Supreme Court, High Court, and District Court Cases,
Central Government Acts of India
5.4M Legal-BERT-base-uncased
InCaseLawBERT
Paul et al. (2023)
Indian Supreme Court, High Court, and District Court Cases,
Central Government Acts of India
5.4M CaseLaw-BERT-base-uncased
CustomInLawBERT
Paul et al. (2023)
Indian Supreme Court, High Court, and District Court Cases,
Central Government Acts of India
5.4M BERT-base-uncased
LexLM
Chalkidis* et al. (2023)
EU Legislation and Case Law, UK Legislation and Case Law,
Canadian Legislation and Case Law, U.S. Case Law and
Contracts, ECHR Case Law, and Indian Case Law
5.8M RoBERTa-base
Legal-XLM-R
Niklaus et al. (2024)
Different Countries Case laws and legislation, US/EU contracts,
and other legal-specific documents
59M XLM-RoBERTa-base
LexT5
T.y.s.s et al. (2024)
EU Legislation and Case Law, UK Legislation and Case Law,
Canadian Legislation and Case Law, U.S. Case Law and
Contracts, ECHR Case Law, and Indian Case Law
5.8M T5-base
Table 3: Key specifications of the evaluated models, including pre-training corpora (with links), document counts, and base models used.
2.2 Tasks and Datasets
Table 2 summarizes key details. Appendix A provides statistics and examples of the datasets which are as follows:
UNFAIR-ToS The UNFAIR-ToS dataset from Chalkidis et al. (2022) is used to identify unfair contractual terms in Terms of Service (ToS) documents from platforms like YouTube, eBay, and Facebook. Each sentence is annotated with one or more of 8 unfairness categories, plus 1 unlabeled class for sentences that do not indicate any potential violation of European consumer law. This makes the task a multi-label classification problem. Labels are based on potential violations of EU consumer protection law . The dataset includes training (5.5k), development (2.3k), and test (1.6k) sets.
LEDGAR The LEDGAR dataset from Chalkidis et al. (2022) is used to classify the principal topic of provisions in Exhibit 10 material contracts (e.g., employment, lease, non-disclosure) filed with the US Securities and Exchange Commission (SEC) via EDGAR . Each provision (paragraph) is labeled with one of 100 contract topics, making it a multi-class classification task. The dataset includes training (60k), development (10k), and test (10k) sets.
LEXDEMOD The LEXDEMOD dataset from Sancheti et al. (2022) detects deontic modality in agent-based contract clauses from lease agreements sourced from the LEDGAR dataset. Each clause (sentence) is annotated with one or more of 6 deontic modality types plus 1 none class , making it a multi-label classification task. Labels are linked to an agent (party) in the sentence, representing their deontic status (e.g., Obligation, Entitlement, Prohibition). The dataset includes training (4.2k), development (330), and test (1.7k) sets. The train/validation/test split is as reported in the original paper.
3 Experiment Setup
We perform task-specific (supervised) fine-tuning using 10 legal-specific LLMs on three datasets: LEDGAR, UNFAIR-ToS, and LEXDEMOD. We consider 10 pre-trained encoder-based legal-specific models for fine-tuning. Nine of these are base-variant encoder models: Legal-BERT Chalkidis et al. (2020), Contracts-BERT Chalkidis et al. (2020), LegalRoBERTa Geng et al. (2021), CaseLaw-BERT Zheng et al. (2021), InLegalBERT, InCaseLawBERT, and CustomInLawBERT Paul et al. (2023), Legal-XLM-R Niklaus et al. (2024), and LexLM Chalkidis* et al. (2023). One large-variant model, PoL-BERT Henderson et al. (2022), is included, as its base version is not present. We also evaluate the encoder-decoder model LexT5 T.y.s.s et al. (2024) (Appendix E), but exclude it from the main results as it is the only model of its kind. Decoder-only models like AdaptLLM Cheng et al. (2024) and SaulLM-7B Colombo et al. (2024) are emerging but custom metrics are not well-supported by the TRL library, which we require in the contract classification case. We leave their benchmarking for future work. A detailed description of each model is provided in Appendix B, and Table 3 summarizes their key characteristics. A detailed experimental setup is provided in Appendix C.
We also compare the 10 legal-specific LLMs with 7 general-purpose LLMs. These include five base variant encoder models: BERT Devlin et al. (2019), RoBERTa-base Liu et al. (2019), DeBERTa He et al. (2021), Longformer Beltagy et al. (2020), and BigBird Zaheer et al. (2020), along with one large variant, RoBERTa-large Liu et al. (2019). Additionally, we compare with the closed-source GPT-3.5-Turbo OpenAI (2022) using zero-shot and one-shot prompting.
UNFAIR-ToS LEDGAR LEXDEMOD
Method Model # Params $\mu$-F1 m-F1 $\mu$-F1 m-F1 $\mu$-F1 m-F1
Zero-shot GPT-3.5-Turbo In Billions 41.4 22.2 70.1 56.7 - -
Few-shot GPT-3.5-Turbo In Billions 64.7 32.5 62.1 51.1 - -
BERT 110M 95.6 81.3 87.6 81.8 - 75.61
RoBERTa-base 125M 95.2 79.2 87.9 82.3 - 75.66
DeBERTa 139M 95.5 80.3 88.2 83.1 - -
Longformer 149M 95.5 80.9 88.2 83.0 - -
BigBird 127M 95.7 81.3 87.8 82.6 - -
Baselines reported
from: Chalkidis (2023),
Chalkidis et al. (2022),
Sancheti et al. (2022)
SFT
(General-
purpose
LLMs)
RoBERTa-large 355M 95.8 81.6 88.6 83.6 - 77.88
Legal-BERT 110M 96.0 82.2 88.2 82.5 81.23 78.01
Contracts-BERT 110M 96.2 83.4 87.9 82.2 80.17 77.71
Legal-RoBERTa 125M 95.4 81.1 87.7 81.9 80.12 76.70
CaseLawBERT 110M 96.1 83.2 87.6 80.9 80.32 77.75
PoL-BERT 340M 94.6 77.9 86.0 79.1 41.35 15.75
InLegalBERT 110M 95.6 81.7 87.9 82.0 80.21 77.89
InCaseLawBERT 110M 95.5 81.1 87.5 82.1 79.16 76.83
CustomInLawBERT 110M 95.5 79.9 87.7 81.8 78.16 75.35
LexLM 124M 95.9 81.7 87.8 81.3 80.39 77.46
Proposed
SFT
(Legal-
specific
LLMs)
Legal-XLM-R 184M 94.9 78.2 87.7 81.7 80.62 77.56
Table 4: Performance of legal-specific and general-purpose LLMs on three tasks: UNFAIR-ToS, LEDGAR, LEXDEMOD. Metrics: micro-F1 ($\mu$-F1) and macro-F1 (m-F1). SFT denotes supervised fine-tuning; zero-shot and few-shot indicate prompting methods. Red highlights best legal-specific, blue highlights best general-purpose performance.
Mean $\pm$ Std
Legal Specific
LLMs
$\mu$-F1 m-F1
Legal-BERT 88.48 ± 6.03 80.90 ± 2.05
Contracts-BERT 88.09 ± 6.55 81.10 ± 2.45
Legal-RoBERTa 87.74 ± 6.24 79.90 ± 2.29
CaseLawBERT 88.01 ± 6.45 80.62 ± 2.23
PoL-BERT 73.98 ± 23.34 57.58 ± 29.58
InLegalBERT 87.90 ± 6.28 80.53 ± 1.87
InCaseLawBERT 87.39 ± 6.67 80.01 ± 2.29
CustomInLawBERT 87.12 ± 7.09 79.02 ± 2.71
LexLM 88.03 ± 6.33 80.15 ± 1.91
Legal-XLM-R 87.74 ± 5.83 79.15 ± 1.82
LexT5 85.60 ± 7.73 76.40 ± 2.66
Table 5: Aggregated scores (Mean $\pm$ Std) across three contract understanding tasks. Red, blue, and green highlights indicate the first, second, and third best performances, respectively.
4 Results and Analysis
Table 4 reports the test results of LLMs across all three tasks, while Table 5 presents the aggregated scores. To address the main research question in Section 1, we run experiments to answer the following Sub-Research Questions (SRQs):
SRQ1: How do legal-specific LLMs perform across different contract understanding tasks compared to general-purpose LLMs?
Table 4 compares legal-specific and general-purpose LLMs. Among general models, RoBERTa-large (355M) performs best overall. However, LLMs such as Contracts-BERT and Legal-BERT (110M) outperform RoBERTa-large on UNFAIR-ToS and LEXDEMOD, respectively, despite having 69% fewer parameters. Other legal LLMs, including CaseLaw-BERT and LexLM, also surpass RoBERTa-large on UNFAIR-ToS. Legal-RoBERTa, CustomInLawBERT, InCaseLawBERT, and InLegalBERT consistently outperform RoBERTa-base, BERT, DeBERTa, and Longformer on UNFAIR-ToS. On LEXDEMOD, Legal-BERT and InLegalBERT again outperform RoBERTa-large. These results highlight that legal-specific base variant LLMs, despite having 64-69% fewer parameters, often outperform larger general-purpose LLMs on domain-specific tasks. RoBERTa-large remains the best model for LEDGAR. Still, Legal-BERT delivers equivalent performance compared to general-purpose base variant models on this task, suggesting that both model size and task characteristics influence performance. The larger legal-specific LLMs may be better suited for LEDGAR. Overall, we conclude that legal-specific base models deliver competitive performance and set new SOTAs on two of the three tasks, demonstrating the effectiveness of domain-specific pretraining, even at the base variant of LLMs.
SRQ2: Which legal-specific LLMs serve as strong baselines for contract understanding?
Table 5 presents aggregated test scores (arithmetic, harmonic, and geometric means) across the three contract understanding tasks. The top three performances are highlighted in red, blue, and green respectively. Despite class imbalance in all tasks, Legal-BERT achieves the highest aggregated $\mu$-F1, while Contracts-BERT leads in m-F1. Across both metrics, the top positions are consistently held by four legal-specific models: Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM. We conclude that these four models, Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM, should be considered strong baselines for contract understanding tasks.
SRQ3: What are the observed limitations of current legal-specific LLMs, and how can these findings guide future legal LLMs development?
Several recent legal-specific LLMs, such as Legal-RoBERTa, CaseLaw-BERT, PoL-BERT, CustomInLawBERT, LexLM, and Legal-XLM-R, are pre-trained on large-scale legal corpora. Models like InLegalBERT and InCaseLawBERT are built on legal-specific base models rather than general-purpose models. However, older legal-specific base-variant LLMs, such as LegalBERT and ContractsBERT, which are pre-trained on just 354k and 76k legal documents respectively (as seen in Table 3), still outperform many recent base-variant legal-specific models (as seen in Table 5). A key limitation of recent legal-specific LLMs is that they are pre-trained on few, or no, diverse contract documents compared to other legal texts like legislation and court cases. We conclude that future legal-specific LLMs should incorporate a more diverse and representative set of contract documents, to improve performance across contract understanding tasks.
5 Conclusion
This study benchmarks 10 legal-specific LLMs against 7 general-purpose LLMs across three contract understanding tasks. Legal-specific base LLMs consistently perform well and set new SOTAs on two tasks despite having fewer parameters. Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM emerge as strong baseline models for contract understanding. However, recent base-variant legal LLMs often underperform due to limited pretraining on diverse contract data. Future work focuses on expanding contract data and evaluating emerging decoder-based legal LLMs.
Limitations
The limited availability of contract benchmark datasets in languages other than English poses a challenge for multilingual extension. Consequently, this study focuses solely on English-language contract tasks, leaving evaluation on non-English data for future work. While encoder-decoder models like LexT5 and decoder-based legal-specific LLMs such as AdaptLLM and SaulLM-7B are emerging, they remain scarce. We therefore defer their benchmarking until more models become available, ensuring fair comparisons. LexT5 is evaluated for exploratory purposes but excluded from the main results. Additionally, this work concentrates on the nuances of contract language and does not assess performance on other legal text types, such as statutes, court decisions, or legal opinions. Future research should extend this evaluation to a broader range of legal genres, acknowledging that no single study can fully capture the entire legal domain.
Ethical Considerations
This study uses only publicly available datasets, LEDGAR, UNFAIR-ToS, and LEXDEMOD, all of which contain contract clauses without personal data. LEDGAR is derived from public U.S. SEC EDGAR filings, UNFAIR-ToS from company Terms of Service, and LEXDEMOD from lease clauses sourced from LEDGAR. This research does not offer legal advice, predict individual outcomes, or automate decisions affecting rights. It focuses solely on evaluating the performance of legal-specific LLMs to inform future tools and research. While these models can support legal professionals, they are not substitutes for legal expertise. We acknowledge potential ethical risks if outputs are misused or inaccurate. By open-sourcing our evaluations, we aim to reduce reliance on proprietary tools, promote transparency, and expand access to legal AI research and development.
Acknowledgment
We utilized commercial AI tools to enhance the clarity, grammar, and style of our manuscript.
Appendix A Dataset Statistics and Illustrative Examples
This appendix presents labeled examples from all three datasets to aid understanding. We provide our own rationales explaining label types, which the datasets do not explicitly include. These explanations clarify why specific labels apply to given clauses (sentences) or provisions (paragraphs), as detailed in Table 6. Legal contract classification involves longer texts than typical NLP tasks like tweets or reviews. Legal-specific Transformer models such as Legal-BERT process up to 512 sub-word tokens, but many LEDGAR paragraphs exceed this limit. Figure 1 shows that numerous LEDGAR paragraphs surpass the standard context window, requiring truncation or other methods to handle long inputs. Additionally, legal texts contain specialized terminology (legalese), increasing classification complexity compared to general text.
Figure 1: Distribution of text lengths, measured in Legal-BERT subword units, across all three datasets
Dataset Example Label
Rationale for Assigned Labels
(Provided by us for better understanding)
This Amendment may be executed by one or more of the
parties hereto on any number of separate counterparts,
and all of said counterparts taken together shall be deemed
to constitute one and the same instrument. This
Amendment may be delivered by facsimile or other
electronic transmission of the relevant signature pages
hereof.
Counterparts
This sentence states that the Amendment may be
executed in multiple counterparts and that together
they form a single agreement, which is a standard
counterparts clause used to validate separately
signed copies as one binding document.
THIS AMENDMENT SHALL BE GOVERNED BY, AND
INTERPRETED IN ACCORDANCE WITH, THE LAW
OF THE STATE OF NEW YORK . The other provisions of
Article IX of the Credit Agreement shall apply to this
Amendment to the same extent as if fully set forth herein.
Governing Laws
This sentence specifies that New York law will
govern and interpret the Amendment, which is a
standard governing law clause that establishes
the legal jurisdiction and framework for resolving
disputes.
LEDGAR
Sublessee leases the Aircraft in its “as is, where is” condition.
The only services, rights, or warranties to which the Sublessee
is entitled to under this Sublease are those to which the
Sublessor is provided under the Prime Lease.
Warranties
The sentence is labeled as warranties because it
defines the rights and guarantees the Sublessee
receives and limits those warranties to what the
Sublessor has under the Prime Lease.
Niantic further reserves the right to remove any User
Content from the Service at any time and without notice
and for any reason.
Content removal
This sentence is labeled as content removal unfair
contractual term because it gives the provider full
control to remove content at any time, for any
reason, and without notice.
amazon reserves the right to refuse service, terminate
accounts , terminate your rights to use amazon services,
remove or edit content , or cancel orders in its sole
discretion.
Unilateral
termination,
Content removal
This sentence is labeled as unilateral termination and
content removal because it allows Amazon to end
services, remove content, or cancel orders at its sole
discretion, without notice, creating an imbalance of
power.
UNFAIR-ToS
Outside the United States and Canada. If you acquired
the if you acquired the application in any other country,
the laws of that country apply.
None
The sentence is labeled as none because it does not
belong to any of the unfair contractual term types
and is actually a fair clause.
[lessee] Lessee will not create or permit to be
created or to remain , and will promptly
discharge , any lien , encumbrance or charge (including
without limitation any mechanic ’s , laborer ’s or
materialman ’s lien ) against the Premises or any part
thereof arising from Lessee ’s actions.
Prohibition,
Obligation
This sentence imposes a prohibition by forbidding
the Lessee (the agent) from creating or allowing liens.
It also imposes an obligation by requiring the Lessee
to promptly remove any such liens.
[tenant] Tenant may, without Landlord’s consent,
before delinquency occurs, contest any such taxes related
to the Personal Property.
Permission
This sentence grants permission to the Tenant (the
agent) to contest taxes without needing the
Landlord’s consent, as long as it’s done before
delinquency.
LEXDEMOD
[landlord] Tenant shall promptly notify Landlord of any
alleged defaults under the CC&Rs and/or the Oil and Gas
Lease .
Entitlement
The Landlord, as the agent, holds an entitlement to
receive notice from the Tenant about alleged defaults.
Table 6: Overview of all three Datasets with Examples, Labels, and Author-Provided Rationales
Appendix B Description of Legal-specific LLMs
Legal-BERT Legal-BERT Chalkidis et al. (2020) is a BERT-base-uncased model (110M parameters) pre-trained on 354K English legal documents, including EU and UK legislation, US contracts, and US and EU court cases. It follows the original BERT pre-training configuration and constructs its sub-word vocabulary from scratch to better capture legal terminology.
Contracts-BERT Contracts-BERT Chalkidis et al. (2020) is a BERT-base-uncased model (110M parameters) pre-trained on 76K US contracts. It follows the original BERT configuration and retains a custom vocabulary tailored to contract language.
Legal-RoBERTa Legal-RoBERTa Geng et al. (2021) builds on the RoBERTa-base model (125M parameters) and continues pre-training on 4.9 GB of legal text, including patent litigation documents, US court cases, and publicly available Google Patents data.
CaseLaw-BERT CaseLaw-BERT Zheng et al. (2021) is a BERT-base-uncased model (110M parameters) pre-trained on 3.4M US federal and state court decisions from the Harvard Case Law corpus. Although originally referred to as Custom Legal-BERT by Zheng et al. (2021), it is later termed CaseLaw-BERT by Chalkidis et al. (2022) to distinguish it from the earlier Legal-BERT of Chalkidis et al. (2020), highlighting its exclusive training on harvard case law. This naming convention is now widely adopted, and we follow the same in this work.
PoL-BERT PoL-BERT Henderson et al. (2022) is a RoBERTa-large model (340M parameters) pre-trained on the Pile-of-Law, a 256GB corpus comprising 10M legal and administrative documents. The dataset spans a wide range of legal domains, including US federal and state court opinions (e.g., CourtListener, SCOTUS filings), regulatory documents (e.g., Federal Register, Code of Federal Regulations, SEC and IRS guidance), legislative texts (e.g., US Bills, US Code, State Codes), and other legal document sources (e.g., ECHR, Eur-Lex, ICJ/PCIJ rulings). It also includes administrative decisions from US agencies (e.g., DOJ, OLC, BVA, NLRB, EOIR, DOL), legal contracts (e.g., EDGAR filings, Atticus contracts, CFPB agreements), educational materials (e.g., open-access casebooks, exam outlines), and publicly available community-driven legal discussions.
InLegalBERT InLegalBERT Paul et al. (2023) builds on Legal-BERT-base-uncased Chalkidis et al. (2020), a legal-specific BERT model (110M parameters) initially pre-trained on 354K English legal documents, including EU and UK legislation, US contracts, and US and EU court cases. It is further pre-trained on 5.4M Indian legal documents, including judgments from the Supreme Court, High Courts, and District Courts, as well as Central Government Acts of India. This extended pre-training enables the model to better capture the linguistic and legal nuances of other jurisdictions, such as Indian jurisprudence.
InCaseLawBERT InCaseLawBERT Paul et al. (2023) builds on CaseLaw-BERT-base-uncased (110M parameters), which is initially pre-trained on 3.4M US federal and state court decisions from the Harvard Case Law corpus. It undergoes further pre-training on 5.4M Indian legal documents, including judgments from the Supreme Court, High Courts, and District Courts, as well as Central Government Acts of India. This additional training enables the model to better capture the linguistic and legal nuances of other jurisdictions, particularly Indian jurisprudence.
CustomInLawBERT CustomInLawBERT Paul et al. (2023) is a BERT-base-uncased model (110M parameters) pre-trained from scratch on 5.4M Indian legal documents, including judgments from the Supreme Court, High Courts, and District Courts, as well as Central Government Acts of India.
LexLMs LexLMs Chalkidis* et al. (2023) include two variants: RoBERTa-base (124M parameters) and RoBERTa-large (340M parameters), both pre-trained from scratch on 5.8M legal documents from multiple English-speaking jurisdictions. The corpus covers a wide range of sources, including EU legislation, EU and ECtHR court decisions, UK legislation and court cases, Indian court decisions, Canadian legislation and court decisions, US court decisions, US legislation, and US contracts. The dataset is designed to ensure broad jurisdictional and document-type coverage, with US legal texts comprising the largest portion. This large-scale, English legal-domain pre-training enables LexLMs to support robust legal language understanding across common law and mixed legal systems.
Legal-XLM-R Legal-XLM-R Niklaus et al. (2024) includes two variants: RoBERTa-base (124M parameters) and RoBERTa-large (340M parameters), both pre-trained from scratch on a multilingual legal corpus comprising 59M documents. The dataset spans 24 languages and five legal text types, including legislation and case law, collected from various jurisdictions such as Germany, Switzerland, the UK, and several other countries. This large-scale, cross-lingual pre-training enables Legal-XLM-R to support legal language understanding across multilingual and multi-jurisdictional contexts.
LexT5 LexT5 T.y.s.s et al. (2024) is a legal-oriented sequence-to-sequence model designed to address the limitations of encoder-only architectures in legal NLP. It is pre-trained on three T5 variants, T5 Small (60M parameters), T5 Base (220M), and T5 Large (770M), using the same 5.8 million legal documents employed for LexLMs Chalkidis* et al. (2023).
UNFAIR-ToS LEDGAR LEXDEMOD
Method Model # Params $\mu$-F1 m-F1 $\mu$-F1 m-F1 $\mu$-F1 m-F1
Legal-BERT 110M 95.20 78.88 88.32 82.10 76.08 72.76
Contracts-BERT 110M 95.13 76.31 87.83 81.97 77.51 72.49
Legal-RoBERTa 125M 94.79 76.16 87.72 81.17 77.01 72.68
CaseLawBERT 110M 95.25 75.24 87.65 81.22 76.71 74.74
PoL-BERT 340M 93.92 68.80 85.76 78.60 48.65 17.43
InLegalBERT 110M 95.43 76.82 87.99 81.54 74.79 71.29
InCaseLawBERT 110M 95.44 75.87 87.41 80.52 78.01 76.75
CustomInLawBERT 110M 95.04 72.44 87.37 80.38 78.26 75.03
LexLM 124M 95.34 77.64 87.84 81.26 78.46 75.61
Proposed
SFT
(Legal-
specific
LLMs)
Legal-XLM-R 184M 94.90 71.89 87.50 81.02 77.59 73.50
Table 7: Validation results for all examined legal-specific LLMs. Model performance is evaluated using micro-F1 ($\mu$-F1) and macro-F1 (m-F1). SFT denotes Supervised Fine-Tuning.
Appendix C Experimental Setup
We use all publicly available legal-specific pre-trained models from Hugging Face. To ensure fair comparison, we adopt the training configuration introduced by Chalkidis et al. (2022) for the LEDGAR and UNFAIR-ToS datasets: a learning rate of $3\text{e-}5$ for all nine encoder-base models and $1\text{e-}5$ for the encoder-large model PoL-BERT Henderson et al. (2022), consistent with the setting used for RoBERTa-large. All models are trained for up to 20 epochs with a batch size of 8, using early stopping with a patience of 3 based on development set performance. For UNFAIR-ToS, we use a maximum sequence length of 128, as in Chalkidis et al. (2022). However, we disable mixed-precision training (i.e., set fp16=False) to ensure stable training, which results in longer training times compared to Chalkidis et al. (2022). For LEDGAR, we reduce the maximum sequence length to 128 (from 512 used in Chalkidis et al. (2022)) to save computational resources and training time. This adjustment is necessary because, as discussed above, fp16 is disabled to ensure stable training, which leads to longer training times, and the LEDGAR dataset contains over 80k sentences, which is large. We observe only a marginal performance drop (0.1-0.4%), which we consider acceptable for efficiency. Each model is trained five times with different random seeds (1-5), and we report the test results of the best seed, following Chalkidis et al. (2022) for a fair baseline comparison. For the LEXDEMOD dataset, we follow the setup proposed by Sancheti et al. (2022), using a learning rate of $2\text{e-}5$ for all encoder-based legal-specific models, including PoL-BERT, consistent with their configuration for RoBERTa-large. We use a batch size of 8 and apply early stopping with a patience of 3. The maximum sequence length is set to 256, as in Sancheti et al. (2022). Each model is trained five times with different random seeds, and we report the average test performance across the three best seeds, following Sancheti et al. (2022) for a fair baseline comparison. We evaluate model performance using micro-F1 ($\mu$-F1) and macro-F1 (m-F1) to account for class imbalance. Additionally, we report the arithmetic mean with standard deviation for micro-F1 ($\mu$-F1) and macro-F1 (m-F1) across tasks. All experiments are conducted on a single NVIDIA V100 GPU. Although the datasets are already publicly available, we will release our code and evaluation scripts to support full reproducibility.
Figure 2: Prompt Templates for Instruction-Based Fine-Tuning of LexT5
Appendix D Additional Result
Table 7 presents the development set results for all examined models across the three datasets. Validation results are reported in the same format as the test set, as detailed in the experimental setup in Appendix C.
Appendix E Other Legal-Specific LLMs Considered
In addition to encoder-based legal-specific LLMs, we experiment with encoder-decoder models, specifically LexT5, for exploratory purposes. Instead of supervised fine-tuning, we adopt instruction-based fine-tuning, which aligns better with encoder-decoder and decoder-only architectures. This approach pairs natural language prompts with clause inputs, enabling the model to generate the appropriate label(s) as output. The instruction templates used are listed in Figure 2. For UNFAIR-ToS, we directly use the prompt from Chalkidis (2023), originally designed for zero-shot prompting. For LEXDEMOD and LEDGAR, we design our own prompts inspired by that style. For our experiments, we use the LexT5 Base model with 220M parameters and apply the same hyperparameter configuration used for encoder-based models, as detailed in Appendix C, to ensure fair comparison. On the UNFAIR-ToS dataset, LexT5 achieves a micro-F1 of 95.4 and a macro-F1 of 79.8. While it does not outperform the encoder-based legal-specific model Contract-BERT, it performs better than the general-purpose baseline RoBERTa-large and surpasses some legal-specific models such as PoL-BERT and Legal-XLM-R. On the LEDGAR dataset, LexT5 achieves a micro-F1 of 84.9 and a macro-F1 of 76.1, which falls short of both general-purpose and legal-specific LLMs. On LEXDEMOD, it attains a micro-F1 of 76.5 and a macro-F1 of 73.3, failing to surpass any general-purpose LLMs and underperforming compared to the majority of encoder-based legal-specific models.
Overall, encoder-decoder legal-specific models such as LexT5 currently underperform compared to encoder-based legal-specific and general-purpose LLMs. However, drawing broad conclusions from a single encoder-decoder model is not fair. While encoder-decoder legal-specific models are beginning to emerge, they remain limited in number. At present, LexT5 T.y.s.s et al. (2024) is the only publicly available model of its kind that we are able to find. Decoder-based legal LLMs, such as AdaptLLM Cheng et al. (2024) and SaulLM 7B Colombo et al. (2024), are introduced in 2024; however, such models remain scarce. As a result, we defer comprehensive benchmarking of encoder-decoder and decoder-only legal-specific LLMs to future work, when a broader range of models becomes available to ensure fair comparison.

### v2

@@
-LLMs for Law: Evaluating Legal-Specific LLMs on Contract Understanding
+Evaluating Customized vs. Generalist Transformer-based Models for Legal Contract Classification
 Abstract
-Despite advances in legal NLP, no comprehensive evaluation covering multiple legal-specific LLMs currently exists for contract classification tasks in contract understanding.
-To address this gap, we present an evaluation of 10 legal-specific LLMs on three English-language contract understanding tasks and compare them with 7 general-purpose LLMs.
-The results show that legal-specific LLMs consistently outperform general-purpose models, especially on tasks requiring nuanced legal understanding.
-Legal-BERT and Contracts-BERT establish new SOTAs on two of the three tasks, despite having 69% fewer parameters than the best-performing general-purpose LLM.
-We also identify CaseLaw-BERT and LexLM as strong additional baselines for contract understanding.
-Our results provide a holistic evaluation of legal-specific LLMs and will facilitate the development of more accurate contract understanding systems.
-LLMs for Law: Evaluating Legal-Specific LLMs on Contract Understanding
-Amrita Singhthanks: These authors contributed equally to this work., H.
-Suhan Karaca11footnotemark: 1, Aditya Joshi, Hye-young Paik, Jiaojiao Jiang School of Computer Science and Engineering University of New South Wales (UNSW), Sydney
+Despite advances in legal NLP, no comprehensive evaluation of Transformer-based models customized for legal tasks (referred to as ‘legal-specific’ models in this paper) exists for contract classification tasks.
+To address this gap, we present an evaluation of 13 legal-specific transformer-based models on 3 English-language contract classification tasks and compare them with 9 generalist models.
+The results show that legal-specific models consistently outperform generalist models, especially on tasks requiring nuanced legal understanding.
+They also help reduce misclassification of rare classes in imbalanced datasets.
+Legal-BERT and Contracts-BERT establish new SOTAs on two of the three tasks, despite having 69% fewer parameters than the best-performing generalist models.
+We also identify CaseLaw-BERT and LexLM as strong additional baselines for contract classification.
+Our results highlight the shortcomings of generalist models, emphasizing the need for domain-specific customization, particularly in the context of legal applications.
+Evaluating Customized vs. Generalist Transformer-based Models for Legal Contract Classification
+Amrita Singh, H.
+Suhan Karaca, Aditya Joshi, Hye-young Paik, Jiaojiao Jiang School of Computer Science and Engineering University of New South Wales (UNSW), Sydney 11footnotetext: These authors contributed equally to this work.
 1 Introduction
-Recent work suggests that open-source legal-specific LLMs offer a promising, cost-effective, and privacy-preserving alternative to general-purpose LLMs Bhambhoria et al.
+Recent work suggests that open-source legal-specific models offer a promising, cost-effective, and privacy-preserving alternative to generalist models Singh et al.
+(2025); Bhambhoria et al.
 (2024); Chalkidis et al.
 (2020).
-Prior Work / Ours
-Legal-Specific LLMs UNFAIR-ToS LEDGAR LEXDEMOD
-Legal-BERT ✓/✓ ✓/✓ ✕/✓
-Contracts-BERT ✕/✓ ✕/✓ ✓/✓
-Legal-RoBERTa ✕/✓ ✕/✓ ✕/✓
-CaseLaw-BERT ✓/✓ ✓/✓ ✕/✓
-PoL-BERT ✕/✓ ✕/✓ ✕/✓
-InLegalBERT ✕/✓ ✕/✓ ✕/✓
-InCaseLawBERT ✕/✓ ✕/✓ ✕/✓
-CustomInLawBERT ✕/✓ ✕/✓ ✕/✓
-LexLM ✕/✓ ✕/✓ ✕/✓
-Legal-XLM-R ✕/✓ ✕/✓ ✕/✓
-Table 1: Comparison of our legal-specific LLMs evaluation and coverage with prior work across three contract datasets (tasks): ‘UNFAIR-ToS’ (Unfair Contractual Terms Identification, Lippi et al.
-(2019); Chalkidis et al.
-(2022)), ‘LEDGAR’ (Contract Provision Topic Classification, Tuggener et al.
-(2020); Chalkidis et al.
-(2022)), and ‘LEXDEMOD’ (Agent-Specific Deontic Modality Detection, Sancheti et al.
-(2022)).
-Terms in inverted commas refer to dataset names, tasks are in parentheses.
-✓= model inclusion, ✕= model exclusion.
-However, despite their advantages, these models remain significantly underutilized in current legal NLP downstream tasks.
-As illustrated in Table 1, legal-specific LLMs are rarely evaluated in prior work on three popular and freely available contract understanding tasks such as Unfair Contractual Terms Identification Lippi et al.
+However, despite their advantages, these models remain significantly underutilized in current legal contract classification tasks, as noted by Singh et al.
+(2025).
+As illustrated in Figure 1, legal-specific models are rarely evaluated, and mostly generalist models are preferred in prior work on three popular and freely available contract classification tasks: Unfair Contractual Terms Identification Lippi et al.
 (2019); Chalkidis et al.
@@
 (2022).
-Despite the legal nature of documents/tasks, researchers have continued to favor general-purpose LLMs over legal-specific LLMs.
-In some cases, legal-specific LLMs are excluded entirely.
+Figure 1: Comparison of our legal-specific models evaluation and coverage with prior work.
+Despite the legal nature of documents/tasks, researchers have continued to favor generalist models over legal-specific models.
+This holds true not only for decoder models (which is the predominant context in which ‘customization’ is used) but also for encoder models that must be adopted to the legal domain.
+In some cases, legal-specific models are excluded entirely.
 For instance, recent studies such as Guha et al.
 (2023) and Singh et al.
-(2024), which explicitly focus on legal downstream tasks, do not include any legal-specific LLMs in their benchmarking evaluations.
-Therefore, this paper addresses the Research Question (RQ): How do legal-specific LLMs perform compared to general-purpose LLMs on nuanced legal tasks like contract understanding?
-Dataset
-Contract
-Type
-Task Task Type
-Train/Dev/
-Test Instances
-Classes
-UNFAIR-ToS
-Chalkidis et al. (2022)
-Terms of
-Service
-(Consumer
-Contract)
-Unfair Contractual
-Terms Identification
-Multi-label
-Classification
-5,532/2,275/
-1,607
-9
-LEDGAR
-Chalkidis et al. (2022)
-Exhibit-10
-Material
-Contract
-Contract Provision
-Topic Classification
-Multi-class
-Classification
-60,000/10,000/
-10,000
-100
-LEXDEMOD
-Sancheti et al. (2022)
-Lease
-Contract
-Agent-Specific Deontic
-Modality Detection
-Multi-label
-Classification
-4,282/330/
-1,777
-7
-Table 2: Overview of Datasets used for Benchmarking Legal-specific LLMs.
-To address this question, we present a comprehensive evaluation of 10 open-source legal-specific LLMs with 7 general-purpose LLMs across the three distinct contract understanding tasks.
-Our results reveal consistent improvements in performance for legal-specific LLMs, particularly on tasks where legal and domain-specific semantics are critical.
+(2024), which explicitly focus on legal downstream tasks, do not include any legal-specific models in their benchmarking evaluations.
+Therefore, this paper addresses the Research Question (RQ): How do legal-specific models perform compared to generalist models on legal contract classification tasks?
+To address this question, we present a comprehensive evaluation of 13 open-source legal-specific models with 9 generalist models across the three distinct contract classification tasks.
+Our results reveal consistent improvements in performance for legal-specific models, particularly on tasks where legal and domain-specific semantics are critical.
 This benchmark serves as a resource for the community, offering a clearer understanding of model suitability and performance across tasks and model types.
-The contributions of this work are as follows: (a) To the best of our knowledge, we present the first benchmarking of multiple legal-specific LLMs across multiple contract understanding tasks; (b) We systematically compare their performance with that of general-purpose LLMs; (c) We identify model strengths, weaknesses, and task-specific challenges, offering insights for future research and deployment.
+The contributions of this work are as follows: (a) To the best of our knowledge, we present the first benchmarking of multiple legal-specific models across multiple contract classification tasks; (b) We systematically compare their performance with that of generalist models; (c) We identify model strengths, weaknesses, and task-specific challenges, offering insights for future research and deployment.
 2 Contract Classification Tasks and Datasets
-2.1 Dataset Selection Desiderata
-Based on following factors, we select the legal contract datasets and tasks:
-Language: English-language contract datasets are selected due to their availability, provide consistent benchmarking for future legal-specific models in the global research community, enable comparison with past benchmarked models.
-Relevance and Diversity: The focus is on contract classification tasks that reflect real-world contract review and analysis challenges, and that test a model’s understanding of legal language, structure, and semantics.
-As shown in Table 2, three distinct tasks are selected, each using a different dataset and representing a unique contract classification scenario in terms of dataset size and number of classes.
-Difficulty: Datasets are chosen where SOTA general-purpose language models do not achieve near-perfect performance Lippi et al.
+To evaluate the effectiveness of models customized for the legal setting vis-a-vis generalist models, we consider the following factors to select the datasets and tasks.
+We select the language to be English due to the availability of datasets and models.
+Publicly available, well-documented datasets are used, each large enough for stable training and evaluation.
+Proprietary, non-public, and very small datasets (under 3K sentences) are avoided to ensure reproducibility and generalizability.
+This criterion modifies and adapts the selection guidelines of Chalkidis et al.
+(2022).
+Datasets are chosen where SOTA generalist language models do not achieve near-perfect performance Lippi et al.
 (2019); Tuggener et al.
@@
 (2022), ensuring that benchmarking legal-specific language models remains challenging.
-Availability & Size: Public, well-documented datasets are used, each large enough for stable training and evaluation.
-Proprietary, non-public, and very small datasets (under 3K sentences) are avoided to ensure reproducibility and generalizability.
-This criterion modifies and adapts the selection guidelines of Chalkidis et al.
+For the chosen tasks to reflect relevance and diversity, we select tasks that test a model’s understanding of legal language, structure, and semantics.
+As shown below, three distinct tasks are selected, each using a different dataset and representing a unique contract classification scenario in terms of dataset size and number of classes.
+Consequently, the following datasets are selected: UNFAIR-ToS: The UNFAIR-ToS dataset (train (5.5k), dev (2.3k), and test (1.6k)) from Chalkidis et al.
+(2022) is a multi-label classification dataset that is used to identify unfair contractual terms in Terms of Service (ToS) documents from online platforms like YouTube.
+Each sentence is annotated with one or more of 8 unfairness categories, plus 1 unlabeled class for sentences that do not indicate any potential violation of European consumer law EU consumer protection law .
+LEDGAR: The LEDGAR dataset (train (60k), dev (10k), and test (10k)) from Chalkidis et al.
+(2022) is used to classify the principal topic of provisions in Exhibit 10 material contracts (e.g., employment, lease, non-disclosure) filed with the US Securities and Exchange Commission (SEC) via EDGAR .
+Each provision (paragraph) is labeled with one of 100 contract topics, making it a multi-class classification task.
+LEXDEMOD: The LEXDEMOD dataset (train (4.2k), dev (330), and test (1.7k)) from Sancheti et al.
+(2022) detects deontic modality in agent-based contract clauses from lease agreements sourced from the LEDGAR dataset.
+Each clause (sentence) is annotated with one or more of 6 deontic modality types plus 1 none class , making it a multi-label classification task.
+Labels are linked to an agent (party) in the sentence, representing their deontic status (e.g., Obligation, Entitlement, Prohibition).
+The train/dev/test split is as reported in the original paper.
+Statistics and examples for all datasets are in Appendix A while the class distribution is in Figure 3.
+UNFAIR-ToS LEDGAR LEXDEMOD
+Method Model # Params $\mu$-F1 m-F1 $\mu$-F1 m-F1 $\mu$-F1 m-F1
+BERT 110M 95.6 81.3 87.6 81.8 - 75.61
+RoBERTa-base 125M 95.2 79.2 87.9 82.3 - 75.66
+DeBERTa 139M 95.5 80.3 88.2 83.1 - -
+Longformer 149M 95.5 80.9 88.2 83.0 - -
+BigBird 127M 95.7 81.3 87.8 82.6 - -
+Baselines reported from: Chalkidis et al. (2022), Sancheti et al. (2022) Generalist Models RoBERTa-large 355M 95.8 81.6 88.6 83.6 - 77.88
+Biswas et al. (2025) Legal-KDD-Student DistilBERT 66M 95.9 76.2 88.2 82.3 - -
+Shin and Kaneko (2025) GATs Hierarchical BERT 110M 95.6 81.5 88.4 83.1 - -
+Llama-3.2 3B 95.9 80.3 85 76.1 76.2 71.4
+Generalist Models Mistral 7B 96.0 80.7 86.4 79.2 76.0 71.2
+Legal-BERT 110M 96.0 82.2 88.2 82.5 81.23 78.01
+Contracts-BERT 110M 96.2 83.4 87.9 82.2 80.17 77.71
+Legal-RoBERTa 125M 95.4 81.1 87.7 81.9 80.12 76.70
+CaseLawBERT 110M 96.1 83.2 87.6 80.9 80.32 77.75
+PoL-BERT 340M 94.6 77.9 86.0 79.1 41.35 15.75
+InLegalBERT 110M 95.6 81.7 87.9 82.0 80.21 77.89
+InCaseLawBERT 110M 95.5 81.1 87.5 82.1 79.16 76.83
+CustomInLawBERT 110M 95.5 79.9 87.7 81.8 78.16 75.35
+LexLM 124M 95.9 81.7 87.8 81.3 80.39 77.46
+Legal-XLM-R 184M 94.9 78.2 87.7 81.7 80.62 77.56
+LexT5 220M 95.4 79.8 84.9 76.1 76.50 73.30
+AdaptLLM 7B 96.5 83.2 85.2 76.8 76.2 70.0
+Proposed Legal-specific Models SaulLM 7B 96.0 81.0 86.6 79.4 76.5 72.7
+Table 1: Performance of legal-specific and generalist models on three tasks: UNFAIR-ToS, LEDGAR, LEXDEMOD.
+Metrics: micro-F1 ($\mu$-F1) and macro-F1 (m-F1).
+Blue highlights the best generalist, and red highlights the best legal-specific model performance.
+Mean $\pm$ Std
+Legal Specific Models $\mu$-F1 m-F1
+Legal-BERT 88.48 ± 6.03 80.90 ± 2.05
+Contracts-BERT 88.09 ± 6.55 81.10 ± 2.45
+Legal-RoBERTa 87.74 ± 6.24 79.90 ± 2.29
+CaseLawBERT 88.01 ± 6.45 80.62 ± 2.23
+PoL-BERT 73.98 ± 23.34 57.58 ± 29.58
+InLegalBERT 87.90 ± 6.28 80.53 ± 1.87
+InCaseLawBERT 87.39 ± 6.67 80.01 ± 2.29
+CustomInLawBERT 87.12 ± 7.09 79.02 ± 2.71
+LexLM 88.03 ± 6.33 80.15 ± 1.91
+Legal-XLM-R 87.74 ± 5.83 79.15 ± 1.82
+LexT5 85.60 ± 7.73 76.40 ± 2.66
+AdaptLLM 85.97 ± 8.31 76.67 ± 5.39
+SaulLM 86.37 ± 7.96 77.70 ± 3.60
+Table 2: Aggregated scores (Mean $\pm$ Std) across three contract classification tasks. Red, blue, and green highlights indicate the first, second, and third best performances, respectively.
+3 Experiment Setup
+We perform task-specific fine-tuning using 13 legal-specific models on three datasets: LEDGAR, UNFAIR-ToS, and LEXDEMOD.
+We consider ten pre-trained encoder-based legal-specific models for fine-tuning.
+Nine of these are base-variant encoder models: Legal-BERT Chalkidis et al.
+(2020), Contracts-BERT Chalkidis et al.
+(2020), LegalRoBERTa Geng et al.
+(2021), CaseLaw-BERT Zheng et al.
+(2021), InLegalBERT, InCaseLawBERT, and CustomInLawBERT Paul et al.
+(2023), Legal-XLM-R Niklaus et al.
+(2024), and LexLM Chalkidis* et al.
 (2023).
+One large-variant model, PoL-BERT Henderson et al.
+(2022), is included, as its base version is not present.
+We also evaluate two legal decoder-based models, AdaptLLM Cheng et al.
+(2024) and SaulLM-7B Colombo et al.
+(2024), along with one legal encoder-decoder model, LexT5 T.y.s.s et al.
+(2024), as it is the only model of its kind among legal-specific decoder-based and encoder-decoder models.
+Table 5 in Appendix B summarizes the key characteristics of each model, and a detailed description of them is provided in the same appendix.
+A detailed experimental setup is provided in Appendix C.
+We compare these 13 legal-specific models with 9 generalist models.
+These include six base variant encoder models: BERT Devlin et al.
+(2019), RoBERTa-base Liu et al.
+(2019), DeBERTa He et al.
+(2021), Longformer Beltagy et al.
+(2020), BigBird Zaheer et al.
+(2020), and DistilBERT Sanh (2019), along with one large variant, RoBERTa-large Liu et al.
+(2019).
+We also compare two generalist decoder-based models, Llama-3.2 Grattafiori et al.
+(2025) and Mistral Jiang et al.
+(2023).
+4 Results and Analysis
+Generalist models: Table 1 reports the test results of models across all three tasks, comparing legal-specific and generalist models.
+Among generalist models, RoBERTa-large (355M) outperforms others, including larger decoder-based models like Llama-3.2 (3B) and Mistral (7B).
+Figure 3 shows that all datasets suffer from extreme long-tail class imbalance.
+Table 1 confirms that decoder-based models perform poorly on rare classes, as reflected in their lower macro-F1 scores, while RoBERTa-large performs consistently better.
+This suggests that encoder-based models are more robust to extreme class imbalance than decoder-based models.
+Despite having fewer parameters, RoBERTa-large achieves higher micro-F1 and macro-F1, demonstrating that bidirectional context learning and discriminative fine-tuning are more crucial than model scale for legal classification tasks with long-tail distributions.
+Legal-specific models: Legal-specific models such as Contracts-BERT and Legal-BERT (110M) outperform RoBERTa-large on UNFAIR-ToS and LEXDEMOD, respectively, despite having 69% fewer parameters.
+RoBERTa-large remains the best model for LEDGAR.
+Still, Legal-BERT delivers equivalent performance compared to generalist base variant models on this task, suggesting that both model size and task characteristics influence performance.
+The larger, legal-specific encoder-based models may be better suited for LEDGAR, as the dataset is very large.
+Overall, legal-specific base models deliver competitive performance and set new SOTAs on two of the three tasks, demonstrating the effectiveness of domain-specific pretraining, even at the base variant of models.
+Generalist models vs.
+Legal-specific models: We further analyze the errors made by the best generalist model (RoBERTa-large) and examine whether the best legal-specific model (Contracts-BERT) corrects them on the Unfair-ToS dataset.
+As shown in Figure 2, RoBERTa-large majorly misclassifies clauses belonging to rare categories, particularly Limitation of Liability and Unilateral Termination, whereas Contracts-BERT correctly classifies them, suggesting that domain-specific pretraining on contract corpora (refer Table 5) enables better representation of rare legal clause types that generalist models fail to capture.
+This analysis is further extended to decoder-based models by comparing generalist and their corresponding legal-specific variants on the LEDGAR dataset, as shown in Figure 2.
+Specifically, Mistral and its legal-specific counterpart SaulLM, both sharing the same base architecture, are compared alongside Llama-3.2 and AdaptLLM, where AdaptLLM is built on Llama-1 but shares the same Llama architecture lineage.
+Across all cases, the results demonstrate that domain-specific pretraining helps in reducing the misclassification of rare classes in long-tail legal text classification.
+Figure 2: Example outputs showing customized models (green squares) correctly predict examples which were mis-classified by generalist models (red squares).
+Legal-specific models as baselines: Table 2 presents aggregated test scores (Mean $\pm$ Std) across the three contract classification tasks.
+Despite class imbalance in all tasks, Legal-BERT achieves the highest aggregated $\mu$-F1, while Contracts-BERT leads in m-F1.
+Across both metrics, the top positions are consistently held by four legal-specific models: Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM.
+We conclude that these four models, Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM, should be considered strong baselines for contract classification tasks.
+Limitations of Recent Legal-Specific Models: Several recent legal-specific models, such as PoL-BERT, CustomInLawBERT, LexLM, Legal-XLM-R are pre-trained on large-scale legal corpora.
+Models like InLegalBERT and InCaseLawBERT are built on legal-specific base models rather than generalist models.
+However, older models like LegalBERT and ContractsBERT, pre-trained on just 354k and 76k legal documents (Table 5), still outperform recent base-variant legal-specific models (Table 2).
+This counterintuitive result can be attributed to a data distribution mismatch: more data only improves performance when it is in-distribution with the task.
+Recent models are trained on a broad mixture of legal genres, including court cases, legislation, and patents, which dilutes the contract-specific signal.
+Consequently, contracts remain underrepresented relative to other genres in these corpora.
+We conclude that future legal-specific models should incorporate a more diverse and representative set of contract documents, balanced alongside other legal genres, to improve performance on contract-based downstream tasks.
+5 Conclusion
+This study benchmarks 13 legal-specific and 9 generalist models across three contract classification tasks.
+Encoder-based models outperform larger decoder-based models in long-tail legal classification, with domain-specific pretraining consistently improving rare class recognition.
+Legal-specific models set new state-of-the-art results on two of the three tasks, despite having 69% fewer parameters than the best-performing generalist model.
+Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM serve as strong baselines for future research.
+Notably, older legal-specific models trained on smaller, more focused corpora outperform recent models trained on larger, genre-diverse corpora, highlighting that in-distribution pretraining is more critical than scale.
+Future work will explore retrieval-augmented approaches with legal-specific models to further enhance performance on rare and underrepresented legal clause types.
+Limitations
+The limited availability of contract benchmark datasets in languages other than English presents a challenge for multilingual extension.
+As a result, this study focuses solely on English-language contract tasks, with evaluation on non-English data left for future research.
+This work also concentrates on the nuances of contract language and does not assess performance on other legal text types, such as statutes, court decisions, or legal opinions.
+Future research should broaden this evaluation to encompass a wider range of legal genres, recognizing that no single study can fully capture the entire legal domain.
+Additionally, this paper focuses on domain generalization alone while selecting the legal domain.
+Ethical Considerations
+This study uses only publicly available datasets, LEDGAR, UNFAIR-ToS, and LEXDEMOD, all of which contain contract clauses from publicly available contract documents.
+LEDGAR is derived from public U.S.
+SEC EDGAR filings, UNFAIR-ToS from company Terms of Service, and LEXDEMOD from lease clauses sourced from LEDGAR.
+This research does not offer legal advice, predict individual outcomes, or automate decisions affecting rights.
+It focuses solely on evaluating the performance of legal-specific models to inform future tools and research.
+While these models can support legal professionals, they are not substitutes for legal expertise.
+We acknowledge potential ethical risks if outputs are misused or inaccurate.
+By open-sourcing our evaluations, we aim to reduce reliance on proprietary tools, promote transparency, and expand access to legal AI research and development.
+Appendix A Dataset Statistics and Illustrative Examples
+Dataset Contract Type Task Task Type Train /Dev/ Test Class
+UNFAIR-ToS Chalkidis et al. (2022) Terms of Service (Consumer Contract) Unfair Contractual Terms Identification Multi-label Classification 5,532/2,275/ 1,607 9
+LEDGAR Chalkidis et al. (2022) Exhibit-10 Material Contract Contract Provision Topic Classification Multi-class Classification 60,000/10,000/ 10,000 100
+LEXDEMOD Sancheti et al. (2022) Lease Contract Agent-Specific Deontic Modality Detection Multi-label Classification 4,282/330/ 1,777 7
+Table 3: Overview of Datasets used for Benchmarking Legal-specific Models.
+This appendix provides an overview of the datasets used for benchmarking legal-specific models in Table 3 and presents labeled examples from all three datasets to aid understanding.
+The class distribution of all three datasets is given in Figure 3.
+We provide our own rationales explaining label types, which the datasets do not explicitly include.
+These explanations clarify why specific labels apply to given clauses (sentences) or provisions (paragraphs), as detailed in Table 4.
+Legal contract classification involves longer texts than typical NLP tasks like tweets or reviews.
+Legal-specific Transformer models such as Legal-BERT process up to 512 sub-word tokens, but many LEDGAR paragraphs exceed this limit.
+Figure 4 shows that numerous LEDGAR paragraphs surpass the standard context window, requiring truncation or other methods to handle long inputs.
+Additionally, legal texts contain specialized terminology (legalese), increasing classification complexity compared to general text.
+Figure 3: Class Distribution across all three datasets
+Figure 4: Distribution of text lengths, measured in Legal-BERT subword units, across all three datasets
+Dataset Example Label Rationale for Assigned Labels (Provided by us for better understanding)
+This Amendment may be executed by one or more of the parties hereto on any number of separate counterparts, and all of said counterparts taken together shall be deemed to constitute one and the same instrument.
+This Amendment may be delivered by facsimile or other electronic transmission of the relevant signature pages hereof.
+Counterparts This sentence states that the Amendment may be executed in multiple counterparts and that together they form a single agreement, which is a standard counterparts clause used to validate separately signed copies as one binding document.
+THIS AMENDMENT SHALL BE GOVERNED BY, AND INTERPRETED IN ACCORDANCE WITH, THE LAW OF THE STATE OF NEW YORK .
+The other provisions of Article IX of the Credit Agreement shall apply to this Amendment to the same extent as if fully set forth herein.
+Governing Laws This sentence specifies that New York law will govern and interpret the Amendment, which is a standard governing law clause that establishes the legal jurisdiction and framework for resolving disputes.
+LEDGAR Sublessee leases the Aircraft in its “as is, where is” condition.
+The only services, rights, or warranties to which the Sublessee is entitled to under this Sublease are those to which the Sublessor is provided under the Prime Lease.
+Warranties The sentence is labeled as warranties because it defines the rights and guarantees the Sublessee receives and limits those warranties to what the Sublessor has under the Prime Lease.
+Niantic further reserves the right to remove any User Content from the Service at any time and without notice and for any reason.
+Content removal This sentence is labeled as content removal unfair contractual term because it gives the provider full control to remove content at any time, for any reason, and without notice.
+amazon reserves the right to refuse service, terminate accounts , terminate your rights to use amazon services, remove or edit content , or cancel orders in its sole discretion.
+Unilateral termination, Content removal This sentence is labeled as unilateral termination and content removal because it allows Amazon to end services, remove content, or cancel orders at its sole discretion, without notice, creating an imbalance of power.
+UNFAIR-ToS Outside the United States and Canada.
+If you acquired the if you acquired the application in any other country, the laws of that country apply.
+None The sentence is labeled as none because it does not belong to any of the unfair contractual term types and is actually a fair clause.
+[lessee] Lessee will not create or permit to be created or to remain , and will promptly discharge , any lien , encumbrance or charge (including without limitation any mechanic ’s , laborer ’s or materialman ’s lien ) against the Premises or any part thereof arising from Lessee ’s actions.
+Prohibition, Obligation This sentence imposes a prohibition by forbidding the Lessee (the agent) from creating or allowing liens.
+It also imposes an obligation by requiring the Lessee to promptly remove any such liens.
+[tenant] Tenant may, without Landlord’s consent, before delinquency occurs, contest any such taxes related to the Personal Property.
+Permission This sentence grants permission to the Tenant (the agent) to contest taxes without needing the Landlord’s consent, as long as it’s done before delinquency.
+LEXDEMOD [landlord] Tenant shall promptly notify Landlord of any alleged defaults under the CC&Rs and/or the Oil and Gas Lease .
+Entitlement The Landlord, as the agent, holds an entitlement to receive notice from the Tenant about alleged defaults.
+Table 4: Overview of all three Datasets with Examples, Labels, and Author-Provided Rationales
+Appendix B Description of Legal-specific Models
 Legal-Specific Model Pre-training Corpora # Doc Base Model
@@
 5.8M T5-base
+AdaptLLM
+Cheng et al. (2024)
+US court opinions from FreeLaw project
+3.6M LLama-1
+SaulLM-7B
+Cheng et al. (2024)
+FreeLaw, English MultiLegal Pile, EDGAR, EuroParl, GovInfo
+Law Stack Exchange, Australian legal corpora, EU and UK
+legislation, court transcripts, and USPTO filings
+- Mistral
 Table 5: Key specifications of the evaluated models, including pre-training corpora (with links), document counts, and base models used.
-2.2 Tasks and Datasets
-Table 2 summarizes key details. Appendix A provides statistics and examples of the datasets which are as follows:
-UNFAIR-ToS The UNFAIR-ToS dataset from Chalkidis et al.
-(2022) is used to identify unfair contractual terms in Terms of Service (ToS) documents from platforms like YouTube, eBay, and Facebook.
-Each sentence is annotated with one or more of 8 unfairness categories, plus 1 unlabeled class for sentences that do not indicate any potential violation of European consumer law.
-This makes the task a multi-label classification problem.
-Labels are based on potential violations of EU consumer protection law .
-The dataset includes training (5.5k), development (2.3k), and test (1.6k) sets.
-LEDGAR The LEDGAR dataset from Chalkidis et al.
-(2022) is used to classify the principal topic of provisions in Exhibit 10 material contracts (e.g., employment, lease, non-disclosure) filed with the US Securities and Exchange Commission (SEC) via EDGAR .
-Each provision (paragraph) is labeled with one of 100 contract topics, making it a multi-class classification task.
-The dataset includes training (60k), development (10k), and test (10k) sets.
-LEXDEMOD The LEXDEMOD dataset from Sancheti et al.
-(2022) detects deontic modality in agent-based contract clauses from lease agreements sourced from the LEDGAR dataset.
-Each clause (sentence) is annotated with one or more of 6 deontic modality types plus 1 none class , making it a multi-label classification task.
-Labels are linked to an agent (party) in the sentence, representing their deontic status (e.g., Obligation, Entitlement, Prohibition).
-The dataset includes training (4.2k), development (330), and test (1.7k) sets.
-The train/validation/test split is as reported in the original paper.
-3 Experiment Setup
-We perform task-specific (supervised) fine-tuning using 10 legal-specific LLMs on three datasets: LEDGAR, UNFAIR-ToS, and LEXDEMOD.
-We consider 10 pre-trained encoder-based legal-specific models for fine-tuning.
-Nine of these are base-variant encoder models: Legal-BERT Chalkidis et al.
-(2020), Contracts-BERT Chalkidis et al.
-(2020), LegalRoBERTa Geng et al.
-(2021), CaseLaw-BERT Zheng et al.
-(2021), InLegalBERT, InCaseLawBERT, and CustomInLawBERT Paul et al.
-(2023), Legal-XLM-R Niklaus et al.
-(2024), and LexLM Chalkidis* et al.
-(2023).
-One large-variant model, PoL-BERT Henderson et al.
-(2022), is included, as its base version is not present.
-We also evaluate the encoder-decoder model LexT5 T.y.s.s et al.
-(2024) (Appendix E), but exclude it from the main results as it is the only model of its kind.
-Decoder-only models like AdaptLLM Cheng et al.
-(2024) and SaulLM-7B Colombo et al.
-(2024) are emerging but custom metrics are not well-supported by the TRL library, which we require in the contract classification case.
-We leave their benchmarking for future work.
-A detailed description of each model is provided in Appendix B, and Table 3 summarizes their key characteristics.
-A detailed experimental setup is provided in Appendix C.
-We also compare the 10 legal-specific LLMs with 7 general-purpose LLMs.
-These include five base variant encoder models: BERT Devlin et al.
-(2019), RoBERTa-base Liu et al.
-(2019), DeBERTa He et al.
-(2021), Longformer Beltagy et al.
-(2020), and BigBird Zaheer et al.
-(2020), along with one large variant, RoBERTa-large Liu et al.
-(2019).
-Additionally, we compare with the closed-source GPT-3.5-Turbo OpenAI (2022) using zero-shot and one-shot prompting.
-UNFAIR-ToS LEDGAR LEXDEMOD
-Method Model # Params $\mu$-F1 m-F1 $\mu$-F1 m-F1 $\mu$-F1 m-F1
-Zero-shot GPT-3.5-Turbo In Billions 41.4 22.2 70.1 56.7 - -
-Few-shot GPT-3.5-Turbo In Billions 64.7 32.5 62.1 51.1 - -
-BERT 110M 95.6 81.3 87.6 81.8 - 75.61
-RoBERTa-base 125M 95.2 79.2 87.9 82.3 - 75.66
-DeBERTa 139M 95.5 80.3 88.2 83.1 - -
-Longformer 149M 95.5 80.9 88.2 83.0 - -
-BigBird 127M 95.7 81.3 87.8 82.6 - -
-Baselines reported
-from: Chalkidis (2023),
-Chalkidis et al. (2022),
-Sancheti et al. (2022)
-SFT
-(General-
-purpose
-LLMs)
-RoBERTa-large 355M 95.8 81.6 88.6 83.6 - 77.88
-Legal-BERT 110M 96.0 82.2 88.2 82.5 81.23 78.01
-Contracts-BERT 110M 96.2 83.4 87.9 82.2 80.17 77.71
-Legal-RoBERTa 125M 95.4 81.1 87.7 81.9 80.12 76.70
-CaseLawBERT 110M 96.1 83.2 87.6 80.9 80.32 77.75
-PoL-BERT 340M 94.6 77.9 86.0 79.1 41.35 15.75
-InLegalBERT 110M 95.6 81.7 87.9 82.0 80.21 77.89
-InCaseLawBERT 110M 95.5 81.1 87.5 82.1 79.16 76.83
-CustomInLawBERT 110M 95.5 79.9 87.7 81.8 78.16 75.35
-LexLM 124M 95.9 81.7 87.8 81.3 80.39 77.46
-Proposed
-SFT
-(Legal-
-specific
-LLMs)
-Legal-XLM-R 184M 94.9 78.2 87.7 81.7 80.62 77.56
-Table 4: Performance of legal-specific and general-purpose LLMs on three tasks: UNFAIR-ToS, LEDGAR, LEXDEMOD.
-Metrics: micro-F1 ($\mu$-F1) and macro-F1 (m-F1).
-SFT denotes supervised fine-tuning; zero-shot and few-shot indicate prompting methods.
-Red highlights best legal-specific, blue highlights best general-purpose performance.
-Mean $\pm$ Std
-Legal Specific
-LLMs
-$\mu$-F1 m-F1
-Legal-BERT 88.48 ± 6.03 80.90 ± 2.05
-Contracts-BERT 88.09 ± 6.55 81.10 ± 2.45
-Legal-RoBERTa 87.74 ± 6.24 79.90 ± 2.29
-CaseLawBERT 88.01 ± 6.45 80.62 ± 2.23
-PoL-BERT 73.98 ± 23.34 57.58 ± 29.58
-InLegalBERT 87.90 ± 6.28 80.53 ± 1.87
-InCaseLawBERT 87.39 ± 6.67 80.01 ± 2.29
-CustomInLawBERT 87.12 ± 7.09 79.02 ± 2.71
-LexLM 88.03 ± 6.33 80.15 ± 1.91
-Legal-XLM-R 87.74 ± 5.83 79.15 ± 1.82
-LexT5 85.60 ± 7.73 76.40 ± 2.66
-Table 5: Aggregated scores (Mean $\pm$ Std) across three contract understanding tasks. Red, blue, and green highlights indicate the first, second, and third best performances, respectively.
-4 Results and Analysis
-Table 4 reports the test results of LLMs across all three tasks, while Table 5 presents the aggregated scores.
-To address the main research question in Section 1, we run experiments to answer the following Sub-Research Questions (SRQs):
-SRQ1: How do legal-specific LLMs perform across different contract understanding tasks compared to general-purpose LLMs?
-Table 4 compares legal-specific and general-purpose LLMs.
-Among general models, RoBERTa-large (355M) performs best overall.
-However, LLMs such as Contracts-BERT and Legal-BERT (110M) outperform RoBERTa-large on UNFAIR-ToS and LEXDEMOD, respectively, despite having 69% fewer parameters.
-Other legal LLMs, including CaseLaw-BERT and LexLM, also surpass RoBERTa-large on UNFAIR-ToS.
-Legal-RoBERTa, CustomInLawBERT, InCaseLawBERT, and InLegalBERT consistently outperform RoBERTa-base, BERT, DeBERTa, and Longformer on UNFAIR-ToS.
-On LEXDEMOD, Legal-BERT and InLegalBERT again outperform RoBERTa-large.
-These results highlight that legal-specific base variant LLMs, despite having 64-69% fewer parameters, often outperform larger general-purpose LLMs on domain-specific tasks.
-RoBERTa-large remains the best model for LEDGAR.
-Still, Legal-BERT delivers equivalent performance compared to general-purpose base variant models on this task, suggesting that both model size and task characteristics influence performance.
-The larger legal-specific LLMs may be better suited for LEDGAR.
-Overall, we conclude that legal-specific base models deliver competitive performance and set new SOTAs on two of the three tasks, demonstrating the effectiveness of domain-specific pretraining, even at the base variant of LLMs.
-SRQ2: Which legal-specific LLMs serve as strong baselines for contract understanding?
-Table 5 presents aggregated test scores (arithmetic, harmonic, and geometric means) across the three contract understanding tasks.
-The top three performances are highlighted in red, blue, and green respectively.
-Despite class imbalance in all tasks, Legal-BERT achieves the highest aggregated $\mu$-F1, while Contracts-BERT leads in m-F1.
-Across both metrics, the top positions are consistently held by four legal-specific models: Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM.
-We conclude that these four models, Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM, should be considered strong baselines for contract understanding tasks.
-SRQ3: What are the observed limitations of current legal-specific LLMs, and how can these findings guide future legal LLMs development?
-Several recent legal-specific LLMs, such as Legal-RoBERTa, CaseLaw-BERT, PoL-BERT, CustomInLawBERT, LexLM, and Legal-XLM-R, are pre-trained on large-scale legal corpora.
-Models like InLegalBERT and InCaseLawBERT are built on legal-specific base models rather than general-purpose models.
-However, older legal-specific base-variant LLMs, such as LegalBERT and ContractsBERT, which are pre-trained on just 354k and 76k legal documents respectively (as seen in Table 3), still outperform many recent base-variant legal-specific models (as seen in Table 5).
-A key limitation of recent legal-specific LLMs is that they are pre-trained on few, or no, diverse contract documents compared to other legal texts like legislation and court cases.
-We conclude that future legal-specific LLMs should incorporate a more diverse and representative set of contract documents, to improve performance across contract understanding tasks.
-5 Conclusion
-This study benchmarks 10 legal-specific LLMs against 7 general-purpose LLMs across three contract understanding tasks.
-Legal-specific base LLMs consistently perform well and set new SOTAs on two tasks despite having fewer parameters.
-Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM emerge as strong baseline models for contract understanding.
-However, recent base-variant legal LLMs often underperform due to limited pretraining on diverse contract data.
-Future work focuses on expanding contract data and evaluating emerging decoder-based legal LLMs.
-Limitations
-The limited availability of contract benchmark datasets in languages other than English poses a challenge for multilingual extension.
-Consequently, this study focuses solely on English-language contract tasks, leaving evaluation on non-English data for future work.
-While encoder-decoder models like LexT5 and decoder-based legal-specific LLMs such as AdaptLLM and SaulLM-7B are emerging, they remain scarce.
-We therefore defer their benchmarking until more models become available, ensuring fair comparisons.
-LexT5 is evaluated for exploratory purposes but excluded from the main results.
-Additionally, this work concentrates on the nuances of contract language and does not assess performance on other legal text types, such as statutes, court decisions, or legal opinions.
-Future research should extend this evaluation to a broader range of legal genres, acknowledging that no single study can fully capture the entire legal domain.
-Ethical Considerations
-This study uses only publicly available datasets, LEDGAR, UNFAIR-ToS, and LEXDEMOD, all of which contain contract clauses without personal data.
-LEDGAR is derived from public U.S.
-SEC EDGAR filings, UNFAIR-ToS from company Terms of Service, and LEXDEMOD from lease clauses sourced from LEDGAR.
-This research does not offer legal advice, predict individual outcomes, or automate decisions affecting rights.
-It focuses solely on evaluating the performance of legal-specific LLMs to inform future tools and research.
-While these models can support legal professionals, they are not substitutes for legal expertise.
-We acknowledge potential ethical risks if outputs are misused or inaccurate.
-By open-sourcing our evaluations, we aim to reduce reliance on proprietary tools, promote transparency, and expand access to legal AI research and development.
-Acknowledgment
-We utilized commercial AI tools to enhance the clarity, grammar, and style of our manuscript.
-Appendix A Dataset Statistics and Illustrative Examples
-This appendix presents labeled examples from all three datasets to aid understanding.
-We provide our own rationales explaining label types, which the datasets do not explicitly include.
-These explanations clarify why specific labels apply to given clauses (sentences) or provisions (paragraphs), as detailed in Table 6.
-Legal contract classification involves longer texts than typical NLP tasks like tweets or reviews.
-Legal-specific Transformer models such as Legal-BERT process up to 512 sub-word tokens, but many LEDGAR paragraphs exceed this limit.
-Figure 1 shows that numerous LEDGAR paragraphs surpass the standard context window, requiring truncation or other methods to handle long inputs.
-Additionally, legal texts contain specialized terminology (legalese), increasing classification complexity compared to general text.
-Figure 1: Distribution of text lengths, measured in Legal-BERT subword units, across all three datasets
-Dataset Example Label
-Rationale for Assigned Labels
-(Provided by us for better understanding)
-This Amendment may be executed by one or more of the
-parties hereto on any number of separate counterparts,
-and all of said counterparts taken together shall be deemed
-to constitute one and the same instrument. This
-Amendment may be delivered by facsimile or other
-electronic transmission of the relevant signature pages
-hereof.
-Counterparts
-This sentence states that the Amendment may be
-executed in multiple counterparts and that together
-they form a single agreement, which is a standard
-counterparts clause used to validate separately
-signed copies as one binding document.
-THIS AMENDMENT SHALL BE GOVERNED BY, AND
-INTERPRETED IN ACCORDANCE WITH, THE LAW
-OF THE STATE OF NEW YORK . The other provisions of
-Article IX of the Credit Agreement shall apply to this
-Amendment to the same extent as if fully set forth herein.
-Governing Laws
-This sentence specifies that New York law will
-govern and interpret the Amendment, which is a
-standard governing law clause that establishes
-the legal jurisdiction and framework for resolving
-disputes.
-LEDGAR
-Sublessee leases the Aircraft in its “as is, where is” condition.
-The only services, rights, or warranties to which the Sublessee
-is entitled to under this Sublease are those to which the
-Sublessor is provided under the Prime Lease.
-Warranties
-The sentence is labeled as warranties because it
-defines the rights and guarantees the Sublessee
-receives and limits those warranties to what the
-Sublessor has under the Prime Lease.
-Niantic further reserves the right to remove any User
-Content from the Service at any time and without notice
-and for any reason.
-Content removal
-This sentence is labeled as content removal unfair
-contractual term because it gives the provider full
-control to remove content at any time, for any
-reason, and without notice.
-amazon reserves the right to refuse service, terminate
-accounts , terminate your rights to use amazon services,
-remove or edit content , or cancel orders in its sole
-discretion.
-Unilateral
-termination,
-Content removal
-This sentence is labeled as unilateral termination and
-content removal because it allows Amazon to end
-services, remove content, or cancel orders at its sole
-discretion, without notice, creating an imbalance of
-power.
-UNFAIR-ToS
-Outside the United States and Canada. If you acquired
-the if you acquired the application in any other country,
-the laws of that country apply.
-None
-The sentence is labeled as none because it does not
-belong to any of the unfair contractual term types
-and is actually a fair clause.
-[lessee] Lessee will not create or permit to be
-created or to remain , and will promptly
-discharge , any lien , encumbrance or charge (including
-without limitation any mechanic ’s , laborer ’s or
-materialman ’s lien ) against the Premises or any part
-thereof arising from Lessee ’s actions.
-Prohibition,
-Obligation
-This sentence imposes a prohibition by forbidding
-the Lessee (the agent) from creating or allowing liens.
-It also imposes an obligation by requiring the Lessee
-to promptly remove any such liens.
-[tenant] Tenant may, without Landlord’s consent,
-before delinquency occurs, contest any such taxes related
-to the Personal Property.
-Permission
-This sentence grants permission to the Tenant (the
-agent) to contest taxes without needing the
-Landlord’s consent, as long as it’s done before
-delinquency.
-LEXDEMOD
-[landlord] Tenant shall promptly notify Landlord of any
-alleged defaults under the CC&Rs and/or the Oil and Gas
-Lease .
-Entitlement
-The Landlord, as the agent, holds an entitlement to
-receive notice from the Tenant about alleged defaults.
-Table 6: Overview of all three Datasets with Examples, Labels, and Author-Provided Rationales
-Appendix B Description of Legal-specific LLMs
+Table 5 summarizes their key characteristic. The detailed description of the legal-specific models used in the experiments is given below:
 Legal-BERT Legal-BERT Chalkidis et al.
@@
 (2023).
-UNFAIR-ToS LEDGAR LEXDEMOD
-Method Model # Params $\mu$-F1 m-F1 $\mu$-F1 m-F1 $\mu$-F1 m-F1
-Legal-BERT 110M 95.20 78.88 88.32 82.10 76.08 72.76
-Contracts-BERT 110M 95.13 76.31 87.83 81.97 77.51 72.49
-Legal-RoBERTa 125M 94.79 76.16 87.72 81.17 77.01 72.68
-CaseLawBERT 110M 95.25 75.24 87.65 81.22 76.71 74.74
-PoL-BERT 340M 93.92 68.80 85.76 78.60 48.65 17.43
-InLegalBERT 110M 95.43 76.82 87.99 81.54 74.79 71.29
-InCaseLawBERT 110M 95.44 75.87 87.41 80.52 78.01 76.75
-CustomInLawBERT 110M 95.04 72.44 87.37 80.38 78.26 75.03
-LexLM 124M 95.34 77.64 87.84 81.26 78.46 75.61
-Proposed
-SFT
-(Legal-
-specific
-LLMs)
-Legal-XLM-R 184M 94.90 71.89 87.50 81.02 77.59 73.50
-Table 7: Validation results for all examined legal-specific LLMs. Model performance is evaluated using micro-F1 ($\mu$-F1) and macro-F1 (m-F1). SFT denotes Supervised Fine-Tuning.
+AdaptLLM AdaptLLM Cheng et al.
+(2024) is a LLaMA-1 model (7B parameters) pre-trained on 3.6M US court opinions from the FreeLaw project Gao et al.
+(2020).
+It converts raw legal texts into reading comprehension tasks, such as summarization, NLI, commonsense reasoning, and text completion, using regex-based patterns, and learns in a self-supervised manner.
+Figure 5: Prompt Templates for Instruction-Based Fine-Tuning of LexT5
+SaulLM-7B SaulLM-7B Colombo et al.
+(2024) is a Mistral-based model (7B parameters) pre-trained on 94B tokens of legal text.
+It combines data from existing sources such as FreeLaw, the English MultiLegal Pile, EDGAR, and EuroParl, along with additional datasets including GovInfo, Law Stack Exchange, Australian legal corpora, EU and UK legislation, court transcripts, and USPTO filings.
+The corpus is curated from both public datasets and web-scraped legal resources to ensure broad legal coverage across jurisdictions.
 Appendix C Experimental Setup
@@
 (2022) for a fair baseline comparison.
-We evaluate model performance using micro-F1 ($\mu$-F1) and macro-F1 (m-F1) to account for class imbalance.
-Additionally, we report the arithmetic mean with standard deviation for micro-F1 ($\mu$-F1) and macro-F1 (m-F1) across tasks.
-All experiments are conducted on a single NVIDIA V100 GPU.
-Although the datasets are already publicly available, we will release our code and evaluation scripts to support full reproducibility.
-Figure 2: Prompt Templates for Instruction-Based Fine-Tuning of LexT5
-Appendix D Additional Result
-Table 7 presents the development set results for all examined models across the three datasets.
-Validation results are reported in the same format as the test set, as detailed in the experimental setup in Appendix C.
-Appendix E Other Legal-Specific LLMs Considered
-In addition to encoder-based legal-specific LLMs, we experiment with encoder-decoder models, specifically LexT5, for exploratory purposes.
-Instead of supervised fine-tuning, we adopt instruction-based fine-tuning, which aligns better with encoder-decoder and decoder-only architectures.
+For decoder-based generalist and legal-specific models, the parameters remain the same.
+However, for the LEDGAR dataset, the number of epochs for models such as Mistral, AdaptLLM, and SaulLM is reduced to 3 instead of 20 due to the long training time (22-24 hours per model).
+This adjustment is made because fp16 is disabled for stable training, and it is necessary for LEDGAR due to its large model size and the resource constraints.
+In addition to encoder-decoder based legal-specific model (LexT5), we adopt instruction-based fine-tuning, which aligns better with encoder-decoder models Wang et al.
+(2023).
 This approach pairs natural language prompts with clause inputs, enabling the model to generate the appropriate label(s) as output.
@@
 For LEXDEMOD and LEDGAR, we design our own prompts inspired by that style.
-For our experiments, we use the LexT5 Base model with 220M parameters and apply the same hyperparameter configuration used for encoder-based models, as detailed in Appendix C, to ensure fair comparison.
-On the UNFAIR-ToS dataset, LexT5 achieves a micro-F1 of 95.4 and a macro-F1 of 79.8.
-While it does not outperform the encoder-based legal-specific model Contract-BERT, it performs better than the general-purpose baseline RoBERTa-large and surpasses some legal-specific models such as PoL-BERT and Legal-XLM-R.
-On the LEDGAR dataset, LexT5 achieves a micro-F1 of 84.9 and a macro-F1 of 76.1, which falls short of both general-purpose and legal-specific LLMs.
-On LEXDEMOD, it attains a micro-F1 of 76.5 and a macro-F1 of 73.3, failing to surpass any general-purpose LLMs and underperforming compared to the majority of encoder-based legal-specific models.
-Overall, encoder-decoder legal-specific models such as LexT5 currently underperform compared to encoder-based legal-specific and general-purpose LLMs.
-However, drawing broad conclusions from a single encoder-decoder model is not fair.
-While encoder-decoder legal-specific models are beginning to emerge, they remain limited in number.
-At present, LexT5 T.y.s.s et al.
-(2024) is the only publicly available model of its kind that we are able to find.
-Decoder-based legal LLMs, such as AdaptLLM Cheng et al.
-(2024) and SaulLM 7B Colombo et al.
-(2024), are introduced in 2024; however, such models remain scarce.
-As a result, we defer comprehensive benchmarking of encoder-decoder and decoder-only legal-specific LLMs to future work, when a broader range of models becomes available to ensure fair comparison.
+We evaluate model performance using micro-F1 ($\mu$-F1) and macro-F1 (m-F1) to account for class imbalance.
+Additionally, we report the arithmetic mean with standard deviation for micro-F1 ($\mu$-F1) and macro-F1 (m-F1) across tasks.
+All experiments are conducted on a single NVIDIA V100 GPU.
+Appendix D Other generalist Models Considered for Exploratory Purposes
+Here, we discuss the closed-source model (GPT-3.5-Turbo) results using prompting techniques, as reported by Chalkidis (2023).
+For UNFAIR-ToS, zero-shot prompting yields micro and macro F1-scores of 41.4 and 22.2, respectively.
+In few-shot prompting, performance improves, achieving micro and macro F1-scores of 64.7 and 32.5, though these scores remain lower than those achieved through legal-specific and generalist model fine-tuning techniques.
+A similar pattern occurs with LEDGAR, where zero-shot prompting results in micro and macro F1-scores of 70.1 and 56.7, respectively, while few-shot prompting produces scores of 62.1 and 51.1.
+Here, performance decreases with few-shot prompting, but task-specific fine-tuning using legal and generalist models remains the preferred approach.
+Thus, task-specific fine-tuning in the legal domain is essential for accurate legal contract classification.